OpenStudy (anonymous):

PLEASE HELP: Identify the vertex and the y-intercept of the graph of the function y = -3(x + 2)2 + 5.

jimthompson5910 (jim_thompson5910):

y = -3(x + 2)^2 + 5 is in the form y = a(x-h)^2 + k where (h,k) is the vertex

jimthompson5910 (jim_thompson5910):

So h = ?? and k = ??

OpenStudy (anonymous):

h=2 k= 5

OpenStudy (anonymous):

THANK YOU

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

y = -3(x + 2)^2 + 5 y = -3(x - (-2))^2 + 5 so h = -2 and k = 5 this means that the vertex is (-2, 5)